the point on x-axis which is equidistant from the point (2,-1,4)(-4,3,10)



Grade:12

1 Answers

bharathteja p
12 years ago

a point on x axis implies y=0 and z=0

so taking the distance formula,

(x-2)2+12+42 = (x+4)2+32+102

so, -4x+4+1+16=8x+16+9+100

solving for x, 12x=104

x=104/12= 26/3

so the point is (26/3, 0, 0)